OpenStudy (anonymous):

A local weather forcaster is accurate 85% of the time when predicting precipitation for the day. What is the probability that she will make correct precipitation predictions 4 days in a row?

OpenStudy (unklerhaukus):

\[P=(85 \% )^ 4\]

OpenStudy (anonymous):

Okay, thanks! I understand it now. Is it 0.85 * 0.85 * 0.85 * 0.85, right?

OpenStudy (unklerhaukus):

yep ~52%

OpenStudy (anonymous):

That's what I just got, too.

OpenStudy (anonymous):

I have another one: how many different ways can you arrange 7 books on a shelf?

OpenStudy (unklerhaukus):

the first book can go in seven possible places, the second can go in 6 possible places the third can go in 5 possible places ...... \[7!=7\times6\times5\times 4\times 3\times 2\times1=5040\]